I have having a problem trying to figure out a new equation for my problem. Here is the set up.
A father is looking for the best option for daycare for his 5 year old son. A home-based option (Option A) charges a flat rate per hour of $15. A center-based option (option B) charges a fixed fee per week of $150 for 40 hours and $10 an hour for any hour over the 40. What is the best option for the father?
The father works 50 hours a week from his office which is located in his home. Both childcare options are equal distance from his house so the driving expenses don’t need to be considered. Both child care options offer a two week vacation time without having to pay child care but still hold the child’s spot in the daycare. This works well with the father who also gets two weeks of vacation time a year.
Here are my two equations.
Option A: (A)y = 15x
Option B1: (B1)y= 150 + 10 (x-40) – day care with 50 hour child care week
But my paper got sent back saying I need to come up with an additional equation for Option B that represents less than 40 hours of daycare is needed to complete the assessment. I am not sure how to set this up. I came up with this.
y=150+10(x-40), x≥40 hours
Is that right??? If so how do prove when the two equations will be equal.
15x=150+10(x-40), x≥40 hours????

Answer by richwmiller(17219)

There is obviously something wrong with the charge of $15 an hour.
We are told the childcare will be needed for 50 hours a week. There is no reason to assess less than 40 hours.
50 hours
y = 15*50=750
y= 150 + 10 (10)=250
less than 40
y=15*30=450

y=150
The plans are equal for ten hours. After ten hours Plan B is better
Plan A is better for less than ten hours but is not an option for the father.